# Boosting Active Defense Persistence: A Two-Stage Defense Framework Combining Interruption and Poisoning Against Deepfake

### v1

@@ IV-D Comparative Analysis of Defense Persistence
This enhanced persistence is attributed to the efficacy of the poisoning component. As detailed in the model discrepancy analysis in Table V, retraining on TSDF-poisoned data fundamentally destroys the deepfake model’s own generative capabilities. For instance, a StarGAN model’s FID score on clean inputs after this process increases from a baseline of 43.77 to 152.93. Since the model’s core ability to generate quality images is compromised, it cannot learn to properly adapt to or bypass the complex interruption perturbation.
This shows the core synergy of our framework. The poisoning component does not merely protect the interruption defense, but turns the attacker’s own retraining attempt into a self-destructive process that destroys their model.
@@ IV-E Visual and Qualitative Analysis
As visualized in Fig. 5, the TSDF method demonstrates broadly effective interference across different attributes, images and models. The outputs generated from images protected by TSDF show significantly more distortion and artifacts compared to those from other SOTA methods, particularly on the AGGAN and StarGAN models. This demonstrates the powerful disruptive capability of the framework. Furthermore, Fig. 7 illustrates the success of the poisoning component. The success of TSDF’s poisoning component is visually confirmed by its strong interference effect on multiple face detectors. While these detectors (S3FD, DSFD, and RetinaFace) correctly identify faces in the original images, they fail completely in protected versions by shifting their bounding boxes, reducing the number of detections, or failing to detect a face.
@@ IV-F Ablation Study
TABLE VIII: Compared with different feature processing strategies, the best result is marked in bold.
@@
To demonstrate the advantage of the Feature Enhancement strategy, we compare it with several alternative feature processing methods, using the Kullback-Leibler (KL) divergence and the Wasserstein distance [29]. As shown in Table VIII, the Feature Enhancement method consistently yields the best interruption performance. For example, for the StarGAN model, the strategy achieves an L2mask score of 0.65. These results are better than those of the KL divergence method, which results in an L2mask of 0.22. The baseline without enhancement performs even lower. This confirms that the Feature Enhancement strategy is more effective in creating powerful and disruptive perturbations. This increased efficiency is crucial, as it leaves sufficient space for the effective integration of the poisoning component.
@@ Conclusion
To address the critical lack of persistence in active defenses, this paper introduces the Two-Stage Defense Framework, a novel approach that synergistically combines interruption and data poisoning. This fusion is achieved through the proposed intensity separation mechanism by integrating both functions into a single perturbation. The method provides a powerful and real-time interruption by distorting forged results at the inference stage. Meanwhile, the poisoning component ensures long-term persistence by interfering with the face detection step in an attacker’s retraining process. This directly overcomes the fundamental vulnerability of conventional defenses, which inevitably fail against such adaptive attacks. Experimental results show that our framework effectively distorts the output of the mainstream deepfake model, while also demonstrating a strong capability to contaminate the face detection results. Ultimately, this work establishes a new paradigm for persistent active defense, moving beyond passively surviving retraining to actively destroying an attacker’s ability to adapt.

### v2

@@ IV-D Analysis of Defense Persistence
This enhanced persistence is attributed to the efficacy of the poisoning component. As detailed in the model discrepancy analysis in Table V, retraining on TSDF data fundamentally destroys the deepfake model’s own generative capabilities. For instance, a StarGAN model’s FID score on clean inputs after this process increases from a baseline of 43.77 to 152.93. Since the model’s core ability to generate quality images is compromised, it cannot learn to properly adapt to or bypass the complex interruption perturbation.
This shows the important synergy of our framework. The poisoning component does not merely protect the interruption defense, but turns the attacker’s own retraining attempt into a self-destructive process that destroys their deepfake model.
@@ IV-E Visual and Qualitative Analysis
As visualized in Fig. 5, the TSDF method demonstrates broadly effective interference across different attributes, images and models. The outputs generated from images protected by TSDF show significantly more distortion and artifacts compared to those from other SOTA methods, particularly on the AGGAN and StarGAN models. This demonstrates the powerful disruptive capability of the framework. Furthermore, Fig. 7 illustrates the success of the poisoning component. The success of TSDF’s poisoning component is visually confirmed by its strong interference effect on multiple face detectors. While these detectors (S3FD, DSFD, and RetinaFace) correctly identify faces in the original images, their performance degrades substantially in protected versions by shifting their bounding boxes, reducing the number of detections or failing to detect a face.
@@ IV-F Ablation Study
TABLE VIII: Compared with different feature processing strategies. The best result is marked in bold.
@@
To isolate and evaluate the contribution of our Feature Enhancement strategy specifically to the interruption component, we conduct a dedicated ablation study. In this experiment, the poisoning module is deactivated across all test configurations to ensure that its effects do not interfere with the comparison of interruption performance. We then compare it with several alternative feature processing methods, using the Kullback-Leibler (KL) divergence and the Wasserstein distance [29]. As shown in Table VIII, the Feature Enhancement method consistently yields the best interruption performance. For example, for the StarGAN model, the strategy achieves an L2mask score of 0.65. These results are better than those of the KL divergence method, which results in an L2mask of 0.22. The baseline without enhancement performs even lower. This confirms that the Feature Enhancement strategy is more effective in creating powerful and disruptive perturbations. These results demonstrate the superior efficacy of our Feature Enhancement strategy, which produces a significantly more disruptive perturbation than the baselines. The high efficacy of this primary interruption is a key requirement for our two-stage paradigm, as it establishes the foundational defense that the poisoning component is designed to make persistent.
@@ Conclusion
To address the critical lack of persistence in active defenses, this paper introduces the Two-Stage Defense Framework, a novel approach that synergistically combines interruption and data poisoning. This fusion is achieved through the proposed intensity separation mechanism by integrating both functions into a single perturbation. The method provides a powerful and real-time interruption by distorting forged results at the inference stage. Meanwhile, the poisoning component ensures long-term persistence by interfering with the face detection step in an attacker’s retraining process. This directly addresses the fundamental vulnerability of conventional defenses, which inevitably fail against such adaptive attacks. Experimental results show that our framework effectively distorts the output of the mainstream deepfake model, while also demonstrating a strong capability to contaminate the face detection results. Ultimately, this work establishes a new paradigm for persistent active defense, moving beyond passively surviving retraining to actively destroying an attacker’s ability to adapt.